1. A business process has a process-capability-ratio Cpk = 0.0. Does this process performance meet the 3-sigma quality control standard? If meet standard, please input number "1", if not, input number "2". Flag this Question Question 41 pts

2. A process fills baby formula into bottles with a target of 3 ounces ± 0.20 ounce. Two hundred bottles of baby formula from the process were sampled. The results showed the average amount of baby formula placed in the bottles to be 3.182 ounces. The standard deviation of the amounts was 0.07 ounce. Please determine the value of Cpk for the process. Numbers only. Keep 3-decimal if not exact, either round up or down is ok.
Answer #1

1) Question 41

Answer is 1.

Cpk is 0.0 when the process average equals either LSL or USL i,e it is at par the specification limit.

2) LSL = 3 - 0.2 = 2.8

USL = 3 + 0.2 = 3.2

3*Sigma = 3*0.07 = 0.21

Cpu = (USL - mean) / (3*Sigma) = (3.2 - 3.182) / 0.21

= 0.086

Cpl = (mean - LSL) / (3*sigma)

= (3.182 - 2.8) / 0.21

= 1.819

Cpk = min[Cpu,Cpl] = min[0.086, 1.819]

= 0.086